•You are 30 years old now and you believe you will be able to save for the next 20 years until you are 50. for 10 years following that and till your retirement at age 60 you will not be able to save because of an increase in your expenses due to kid’s college, wedding, etc.

•If you want to guarantee yourself $100,000 per year starting on your 61th birthday how much should you save each year for the next 20 years. Starting at the end of this year.

•Assume that your investment is going to yield 8% and you are likely to live until you are 80.

PLEASE SHOW WORK USING FINANCIAL CALCULATOR. EXAMPLE: N=5, I/Y = 5
Answer #1
PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 100000*((1-(1+ 8/100)^-20)/(8/100))
PV = 981814.74
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =100000
I/Y =8
N =20
FV = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(8/(100),20,,PV,)
Future value = present value*(1+ rate)^time
981814.74 = Present value*(1+0.08)^10
Present value = 454770.19
Using Calculator: press buttons "2ND"+"FV" then assign
FV =-981814.74
I/Y =8
N =10
PMT = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(0.08,10,,981814.74,)
FVOrdinary Annuity = C*(((1 + i/100)^n -1)/(i/100))
C = Cash flow per period
i = interest rate
n = number of payments
454770.19= Cash Flow*(((1+ 8/100)^20-1)/(8/100))
Cash Flow = 9937.73
Using Calculator: press buttons "2ND"+"FV" then assign
FV =454770.19
I/Y =8
N =20
PV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(8/(100),20,,454770.19,)